Laboratory Assignments

  1. A rock dropped off the roof and reaches the ground in 2.6 seconds. What was the velocity at impact?   

                                                                                                                                                       [2 marks]                                                                                           

  1. How long will it take an object that falls from rest to fall a distance of 100 m?                           [2 marks]

  1. As ball falls towards ground, its velocity                                                                                        [2 mark]

  1. Increases
  2. Decreases
  3. Remains constant
  4. Becomes zero

  1. Acceleration of free fall depends on the                                                                                         [2 mark]

  1. Surface
  2. Weight of object
  3. Distance from centre of earth
  4. Size of object
Answer #1

1) using the equation of linear motion we have

v = u + gt

v = 0 + 9.8*2.6

= 25.48 m/s

2) S = ut + 0.5at^2

100 = 0 + 0.5*9.8*t^2

t = 4.52 s

3) the velocity of it will increase

so option a) increase is the correct answer

4) acceleration due to gravity changes with height

so option c) distance from centre of earth is the correct answer